# Maël Dumas LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France Anthony Perez LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France Mathis Rocton Algorithms and Complexity Group, TU Wien, Vienna, Austria Ioan Todinca LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France

### v3

@@ 1 Introduction
Parameterized algorithms are among the most natural approaches to tackle NP-hard optimization problems [14]. In particular, they have been very successful in dealing with so-called edge modification problems on graphs: given as input an arbitrary graph $G=(V,E)$ and a parameter $k\in\mathbb{N}$, the goal is to transform $G$ into a graph with some specific properties (i.e., belonging to a specific graph class $\mathcal{G}$) by adding and/or deleting at most $k$ edges. Parameterized algorithms (also called FPT for fixed parameter tractable) aim at a time complexity of type $f(k)\cdot n^{O(1)}$, where $f$ is some computable function, hence the combinatorial explosion is restricted to parameter $k$.
@@
When the target class $\mathcal{G}$ is characterized by a finite family of forbidden induced subgraphs, modification problems are FPT by a result of Cai [9]. Indeed, as long as the graph contains one of the forbidden subgraphs, one can try each possibility to correct this obstruction and branch by recursive calls. On each branch, the budget $k$ is strictly diminished, therefore the whole algorithm has a number of calls bounded by some function $f(k)$. The situation is more complicated when the target class $\mathcal{G}$ is characterized by an infinite family of forbidden induced subgraphs. Nonetheless, a large literature is devoted to edge modification problems towards chordal graphs (where we forbid all induced cycles with at least four vertices) as well as sub-classes of chordal graphs, typically obtained by requiring some fixed set of obstructions, besides the long cycles. Observe that, in this case, the situation remains relatively simple if we restrict ourselves to edge completion problems, where we are only allowed to add edges to the input graphs. Indeed, in this case, if a graph has an induced cycle of length longer than $k+3$, it cannot be made chordal by adding at most $k$ edges. Therefore we can use again the approach of Cai to deal with cycles of length at most $k+3$ and other obstructions, and either the algorithm finds a solution in $f(k)$ recursive calls, or we can conclude that we have a no-instance. The cases of edge deletion problems (where we are only allowed to remove edges) and edge editing problems (where we are allowed to both remove edges or add missing edges) are more complicated, since even long cycles can be eliminated by a single edge removal. Therefore more efforts and more sophisticated techniques were necessary in these situations, but several such problems turned out to be FPT [11, 16, 17]. The interested reader can refer to [12] for a broad and comprehensive survey on parameterized algorithms for edge modification problems.
@@
We focus here on a sub-family of parameterized algorithms, namely on kernelization. The goal of kernelization is to provide a polynomial algorithm transforming any instance $(I,k)$ of the problem into an equivalent instance $(I^{\prime},k^{\prime})$ where $k^{\prime}$ is upper bounded by some function of $k$ (in our case we will simply have $k^{\prime}\leqslant k$), and the size of the new instance $I^{\prime}$ is upper bounded by some function $g(k)$. Hence the size of the reduced instance does not depend on the size of the original instance. Kernels are obtained through a set of reduction rules. While kernelization is possible for all FPT problems (the two notions are actually equivalent), the interesting question is whether a given FPT problem admits polynomial kernels, where the size of the reduced instance is bounded by some polynomial in $k$. Note that, under some complexity assumptions, not all FPT problems admit polynomial kernels [5, 6, 7, 10, 25, 33].
@@
In this paper we focus on Block Graph Editing, Block Graph Deletion and all three variants of modification problems towards strictly chordal graphs. Block graphs are a well-studied subclass of chordal graphs, that are diamond-free (Figure 2). They are moreover the graphs in which every biconnected component induces a clique. Notice that this observation allows for a simple polynomial-time algorithm for Block Graph Completion since one needs to turn every biconnected component of the input graph into a clique. Strictly chordal graphs are another subclass of chordal graphs, also known as block duplicate graphs [30, 31, 24]. They can be obtained from block graphs by repeatedly choosing some cut-vertex $u$ and adding a true twin $v$ of $u$, that is a vertex $v$ adjacent to $u$ and all neighbors of $u$ [24]. They can also be characterized as dart, gem-free chordal graphs (see Figure 2 and next section) and are thus ptolemaic (i.e. chordal distance-hereditary [8]). Strictly chordal graphs are also known to be a subclass of 4-leaf power graphs [31], and a super-class of 3-leaf power graphs [16].
@@
Kernelization for chordal completion goes back to the ’90s and the seminal paper of Kaplan, Shamir and Tarjan [29]. In this work, the authors provide an $O(k^{3})$ vertex-kernel for Chordal Completion (also known as Minimum Fill-In) that was later improved to $O(k^{2})$ by a tighter analysis [37]. Since then, several authors addressed completion, deletion and/or editing problems towards sub-classes of chordal graphs, such as 3-leaf power graphs [3], split and threshold graphs [26], proper interval graphs [4], trivially perfect graphs [2, 19, 22, 26, 20] or ptolemaic graphs [13]. We provide a general picture for some of those classes Figure 1.
@@
chordal [29, 37]$4$-leaf powerintervalptolemaic [13]trivially perfect [2, 20]strictly chordalsplit [2, 27]threshold [18]proper interval [4]$3$-leaf power [3]blockpoly-kernel $\downarrow$ CompletionEditing, Deletion
@@
Figure 1: Kernelization status of subclasses of chordal graphs. A class below the red or blue line indicates that the corresponding completion, deletion or editing problem admits a polynomial kernel. All the edge modification problems for the presented classes are NP-complete at the exception of the ones for $4$-leaf power for which the complexity is unknown, Split Editing which is surprisingly in P [27], and Block Graph Completion which is also in $P$ as observed previously.
@@
All these classes have in common that they can be defined as chordal graphs, plus a constant number of obstructions. Moreover, several known results deal with classes that share a very common feature (see e.g. [4, 32, 3, 20]). Very informally, the target class $\mathcal{G}$ admits a tree-like decomposition, in the sense that the vertices of any graph $H\in{\mathcal{G}}$ can be partitioned into clique modules (having the same neighborhood in $H$), and these modules can be mapped onto the nodes of a decomposition tree, the structure of the tree describing the adjacencies between modules. Therefore, if an arbitrary graph $G$ can be transformed into graph $H$ by at most $k$ edge additions or deletions, at most $2k$ modules can be affected by the modifications. By removing the affected nodes from the decomposition tree, we are left with several components (chunks) that correspond, in the initial graph $G$ as well as in $H$, to induced subgraphs that may be large but that already belong to the target class. Moreover, these chunks are attached to the rest of graph $G$ in a very regular way, through one or two nodes of the decomposition tree. The kernelization algorithms need to analyze these chunks and provide reduction rules, typically by ensuring a small number of nodes in the decomposition tree, plus the fact that each node corresponds to a module of small size. However, finding a general framework that captures this particular decomposition does not seem to be an easy task as each class exhibits particular structural properties that may need to be dealt with independently.
@@
As mentioned earlier, when restricted to completion problems any input graph cannot have large induced cycles and this property may be exploited to design kernelization algorithms. This observation lead to the following conjecture by Bessy and Perez [4].
Conjecture 1 ([4]).
@@
Several other questions remain open, for example it is not known whether chordal deletion or chordal editing admit polynomial kernels [12].
@@ 2 Preliminaries
Definition 2 (Critical clique graph).
@@
Block graphs are graphs in which every biconnected component is a clique. They can also be characterized as chordal graphs that do not contain diamonds as induced subgraph [1] (see Figure 2). A natural generalization of block graphs are strictly chordal graphs, also known as block duplicate graphs, that are obtained from block graphs by adding true twins [30, 31, 24]. Note that by definition the critical clique graph $\mathcal{C}(G)$ is a block graph.
Theorem 4 (Theorem 8 [35]).
@@
Any minimal separator $S$ of chordal graph $G$ is the intersection of two maximal cliques $\Omega,\Omega^{\prime}$ of $G$, and moreover $S$ is a minimal $ab$-separator for any $a\in\Omega\setminus S$, $b\in\Omega^{\prime}\setminus S$ — in particular, $S$ is strictly included in each of the cliques (see, e.g., Theorem 7 and Lemma 5 in [23]). By construction of $G$, any of its maximal cliques is either a maximal clique of $G_{1}$ or of $G_{2}$, or it is exactly $S_{1}\cup S_{2}$.
@@
We are left with the case when $\Omega=S_{1}\cup S_{2}$ and $\Omega^{\prime}$ is contained in one of the two graphs $G_{i}$, say in $G_{1}$. Therefore $S\subseteq S_{1}$. If this inclusion is strict, we claim that $S$ is also a minimal separator of $G_{1}$. Indeed $S$ is a minimal separator in $G$ for any $a\in\Omega\setminus S$ and $b\in\Omega^{\prime}\setminus S$, therefore we can choose $a\in S_{1}\setminus S$. Again $S$ separates $a$ and $b$ in the subgraph $G_{1}$ of $G$, and since $S$ is in the common neighborhood of $a$ and $b$ in $G_{1}$, it is necessarily a minimal separator of $G$. We deduce that, if $S$ is not a minimal separator of $G_{1}$, we must have $S=S_{1}$. We conclude that $S$ is strictly contained in the maximal clique $\Omega^{\prime}$ of $G_{1}$, which proves our lemma. ∎
@@
Proof.
Observe that minimal separators of block graphs are single vertices since the biconnected components of block graphs are cliques. From Lemma 2, $G$ is chordal and its minimal separators are the ones of $G_{1}$ and $G_{2}$ plus possibly $S_{1}$ (resp. $S_{2}$) if it is a not maximal clique of $G_{1}$ (resp. $G_{2}$). Hence minimal separators of $G$ are single vertices and it follows that $G$ is a block graph. ∎
Lemma 4.
@@
Proof.
From Lemma 2, $G$ is chordal and its minimal separators are the ones of $G_{1}$ and $G_{2}$ plus possibly $S_{1}$ or $S_{2}$ if they are not maximal cliques. If $S_{i}$ intersects exactly one maximal clique, $i\in\{1,2\}$, it is clear that it does not intersect any minimal separator of $G_{i}$. If $S_{i}$ is a critical clique, we claim that either $S_{i}$ is a minimal separator of $G_{i}$ or does not intersect any minimal separator of $G_{i}$. Indeed, if there is a minimal (clique) separator $S$ of $G_{i}$ such that $S_{i}\subset S$, then since $S_{i}$ is a critical clique, there exist $u\in S_{i},v\in S\setminus S_{i}$ and $w\in V(G_{i})\setminus S$ such that $w$ is adjacent to exactly one of $u$ and $v$. Suppose w.l.o.g. that $w$ is adjacent to $u$ and not to $v$, then $u$ is in a minimal $vw$-separator that intersects $S$ and is not equal to it. This is a contradiction since $G_{i}$ is strictly chordal and from Theorem 4 its minimal separators are pairwise disjoint. It follows that the minimal separators of $G$ are pairwise disjoint and by Theorem 4 $G$ is strictly chordal. ∎
@@
Observation 1.
Let $G_{1}=(V_{1},E_{1}),\dots,G_{r}=(V_{r},E_{r})$ be disjoint strictly chordal graphs and for each $i\in\{1,\dots,r\}$ let $S_{i}\subseteq V_{i}$ be a critical clique, a maximal clique or a set that intersects exactly one maximal clique of $G_{i}$. If $S_{1}$ is a critical clique or intersects exactly one maximal clique of $G_{1}$, then $S_{1}$ is a critical clique or intersects exactly one maximal clique of $(G_{1},S_{1})\otimes(G_{2},S_{2})$. This implies by Lemma 4 that $(G_{1},S_{1})\otimes(\bigcup_{2\leqslant i\leqslant r}G_{i},\bigcup_{2\leqslant i\leqslant r}S_{i})$ is strictly chordal (see Figure 3).
@@
We are left with Item iii, which can be proven in a similar way as the previous one. Using the same notation, the only vertices of $T_{A^{\prime}}$ whose degree may increase in $T_{A}$ are in $N$. Therefore, every connected component of $T_{A^{\prime}}\setminus(A^{\prime}\cup d(T_{A^{\prime}}))$ that does not contain vertices of $N$ is also a connected component of $T_{A}\setminus(A\cup d(T_{A}))$. Moreover, a connected component of $T_{A^{\prime}}\setminus(A^{\prime}\cup d(T_{A^{\prime}}))$ may contain vertices of $N$ only if these vertices have degree at most $2$ in $T_{A^{\prime}}$ and are not in $A$. In particular $|N|\leqslant 2$and both vertices lie in the same connected component of $T_{A^{\prime}}\setminus(A^{\prime}\cup d(T_{A^{\prime}}))$ (if any). Therefore, this connected component may be split in two disjoint ones in $T_{A}\setminus(A\cup d(T_{A}))$ if the vertices of $N$ are removed. Moreover, an additional connected component corresponding to the path between $a$ and $u$ may exist in $T_{A}$. Hence, we can only increase the number of connected components of $T_{A}\setminus(A\cup d(T_{A}))$ by compared to $T_{A^{\prime}}\setminus(A^{\prime}\cup d(T_{A^{\prime}}))$. Since the property is verified for $|A|=1$, it follows by induction that there are $2\cdot|A|$ connected components in $T_{A}\setminus(A\cup d(T_{A}))$. ∎

### v4

@@ Abstract
keywords:
parameterized complexity, kernelization, graph modification, block graphs, strictly chordal graphs
@@ 1 Introduction
Parameterized algorithms are among the most natural approaches to tackle NP-hard optimization problems (14). In particular, they have been very successful in dealing with so-called edge modification problems on graphs: given as input an arbitrary graph $G=(V,E)$ and a parameter $k\in\mathbb{N}$, the goal is to transform $G$ into a graph with some specific properties (i.e., belonging to a specific graph class $\mathcal{G}$) by adding and/or deleting at most $k$ edges. Parameterized algorithms (also called FPT for fixed parameter tractable) aim at a time complexity of type $f(k)\cdot n^{O(1)}$, where $f$ is some computable function, hence the combinatorial explosion is restricted to parameter $k$.
@@
When the target class $\mathcal{G}$ is characterized by a finite family of forbidden induced subgraphs, modification problems are FPT by a result of Cai (9). Indeed, as long as the graph contains one of the forbidden subgraphs, one can try each possibility to correct this obstruction and branch by recursive calls. On each branch, the budget $k$ is strictly diminished, therefore the whole algorithm has a number of calls bounded by some function $f(k)$. The situation is more complicated when the target class $\mathcal{G}$ is characterized by an infinite family of forbidden induced subgraphs. Nonetheless, a large literature is devoted to edge modification problems towards chordal graphs (where we forbid all induced cycles with at least four vertices) as well as sub-classes of chordal graphs, typically obtained by requiring some fixed set of obstructions, besides the long cycles. Observe that, in this case, the situation remains relatively simple if we restrict ourselves to edge completion problems, where we are only allowed to add edges to the input graphs. Indeed, in this case, if a graph has an induced cycle of length longer than $k+3$, it cannot be made chordal by adding at most $k$ edges. Therefore we can use again the approach of Cai to deal with cycles of length at most $k+3$ and other obstructions, and either the algorithm finds a solution in $f(k)$ recursive calls, or we can conclude that we have a no-instance. The cases of edge deletion problems (where we are only allowed to remove edges) and edge editing problems (where we are allowed to both remove edges or add missing edges) are more complicated, since even long cycles can be eliminated by a single edge removal. Therefore more efforts and more sophisticated techniques were necessary in these situations, but several such problems turned out to be FPT (11, 16, 17). The interested reader can refer to (12) for a broad and comprehensive survey on parameterized algorithms for edge modification problems.
@@
We focus here on a sub-family of parameterized algorithms, namely on kernelization. The goal of kernelization is to provide a polynomial algorithm transforming any instance $(I,k)$ of the problem into an equivalent instance $(I^{\prime},k^{\prime})$ where $k^{\prime}$ is upper bounded by some function of $k$ (in our case we will simply have $k^{\prime}\leqslant k$), and the size of the new instance $I^{\prime}$ is upper bounded by some function $g(k)$. Hence the size of the reduced instance does not depend on the size of the original instance. Kernels are obtained through a set of reduction rules. While kernelization is possible for all FPT problems (the two notions are actually equivalent), the interesting question is whether a given FPT problem admits polynomial kernels, where the size of the reduced instance is bounded by some polynomial in $k$. Note that, under some complexity assumptions, not all FPT problems admit polynomial kernels (5, 6, 7, 10, 25, 33).
@@
In this paper we focus on Block Graph Editing, Block Graph Deletion and all three variants of modification problems towards strictly chordal graphs. Block graphs are a well-studied subclass of chordal graphs, that are diamond-free (Figure 2). They are moreover the graphs in which every biconnected component induces a clique. Notice that this observation allows for a simple polynomial-time algorithm for Block Graph Completion since one needs to turn every biconnected component of the input graph into a clique. Strictly chordal graphs are another subclass of chordal graphs, also known as block duplicate graphs (30, 31, 24). They can be obtained from block graphs by repeatedly choosing some cut-vertex $u$ and adding a true twin $v$ of $u$, that is a vertex $v$ adjacent to $u$ and all neighbors of $u$ (24). They can also be characterized as dart, gem-free chordal graphs (see Figure 2 and next section) and are thus ptolemaic (i.e. chordal distance-hereditary (8)). Strictly chordal graphs are also known to be a subclass of 4-leaf power graphs (31), and a super-class of 3-leaf power graphs (16).
@@
Kernelization for chordal completion goes back to the ’90s and the seminal paper of Kaplan, Shamir and Tarjan (29). In this work, the authors provide an $O(k^{3})$ vertex-kernel for Chordal Completion (also known as Minimum Fill-In) that was later improved to $O(k^{2})$ by a tighter analysis (37). Since then, several authors addressed completion, deletion and/or editing problems towards sub-classes of chordal graphs, such as 3-leaf power graphs (3), split and threshold graphs (26), proper interval graphs (4), trivially perfect graphs (2, 19, 22, 26, 20) or ptolemaic graphs (13). We provide a general picture for some of those classes Figure 1.
@@
chordal (29, 37)$4$-leaf powerintervalptolemaic (13)trivially perfect (2, 20)strictly chordalsplit (2, 27)threshold (18)proper interval (4)$3$-leaf power (3)blockpoly-kernel $\downarrow$ CompletionEditing, Deletion
@@
Figure 1: Kernelization status of subclasses of chordal graphs. A class below the red or blue line indicates that the corresponding completion, deletion or editing problem admits a polynomial kernel. All the edge modification problems for the presented classes are NP-complete at the exception of the ones for $4$-leaf power for which the complexity is unknown, Split Editing which is surprisingly in P (27), and Block Graph Completion which is also in $P$ as observed previously.
@@
All these classes have in common that they can be defined as chordal graphs, plus a constant number of obstructions. Moreover, several known results deal with classes that share a very common feature (see e.g. (4, 32, 3, 20)). Very informally, the target class $\mathcal{G}$ admits a tree-like decomposition, in the sense that the vertices of any graph $H\in{\mathcal{G}}$ can be partitioned into clique modules (having the same neighborhood in $H$), and these modules can be mapped onto the nodes of a decomposition tree, the structure of the tree describing the adjacencies between modules. Therefore, if an arbitrary graph $G$ can be transformed into graph $H$ by at most $k$ edge additions or deletions, at most $2k$ modules can be affected by the modifications. By removing the affected nodes from the decomposition tree, we are left with several components (chunks) that correspond, in the initial graph $G$ as well as in $H$, to induced subgraphs that may be large but that already belong to the target class. Moreover, these chunks are attached to the rest of graph $G$ in a very regular way, through one or two nodes of the decomposition tree. The kernelization algorithms need to analyze these chunks and provide reduction rules, typically by ensuring a small number of nodes in the decomposition tree, plus the fact that each node corresponds to a module of small size. However, finding a general framework that captures this particular decomposition does not seem to be an easy task as each class exhibits particular structural properties that may need to be dealt with independently.
@@
As mentioned earlier, when restricted to completion problems any input graph cannot have large induced cycles and this property may be exploited to design kernelization algorithms. This observation lead to the following conjecture by Bessy and Perez (4).
Conjecture 1 ((4))
@@
Several other questions remain open, for example it is not known whether chordal deletion or chordal editing admit polynomial kernels (12).
@@ 2 Preliminaries
Definition 2 (Critical clique graph)
@@
Block graphs are graphs in which every biconnected component is a clique. They can also be characterized as chordal graphs that do not contain diamonds as induced subgraph (1) (see Figure 2). A natural generalization of block graphs are strictly chordal graphs, also known as block duplicate graphs, that are obtained from block graphs by adding true twins (30, 31, 24). Note that by definition the critical clique graph $\mathcal{C}(G)$ is a block graph.
Theorem 4 (Theorem 8 (35))
@@
Any minimal separator $S$ of chordal graph $G$ is the intersection of two maximal cliques $\Omega,\Omega^{\prime}$ of $G$, and moreover $S$ is a minimal $ab$-separator for any $a\in\Omega\setminus S$, $b\in\Omega^{\prime}\setminus S$ — in particular, $S$ is strictly included in each of the cliques (see, e.g., Theorem 7 and Lemma 5 in (23)). By construction of $G$, any of its maximal cliques is either a maximal clique of $G_{1}$ or of $G_{2}$, or it is exactly $S_{1}\cup S_{2}$.
@@
We are left with the case when $\Omega=S_{1}\cup S_{2}$ and $\Omega^{\prime}$ is contained in one of the two graphs $G_{i}$, say in $G_{1}$. Therefore $S\subseteq S_{1}$. If this inclusion is strict, we claim that $S$ is also a minimal separator of $G_{1}$. Indeed $S$ is a minimal separator in $G$ for any $a\in\Omega\setminus S$ and $b\in\Omega^{\prime}\setminus S$, therefore we can choose $a\in S_{1}\setminus S$. Again $S$ separates $a$ and $b$ in the subgraph $G_{1}$ of $G$, and since $S$ is in the common neighborhood of $a$ and $b$ in $G_{1}$, it is necessarily a minimal separator of $G$. We deduce that, if $S$ is not a minimal separator of $G_{1}$, we must have $S=S_{1}$. We conclude that $S$ is strictly contained in the maximal clique $\Omega^{\prime}$ of $G_{1}$, which proves our lemma.
@@
Proof 2.7.
Observe that minimal separators of block graphs are single vertices since the biconnected components of block graphs are cliques. From Lemma 2, $G$ is chordal and its minimal separators are the ones of $G_{1}$ and $G_{2}$ plus possibly $S_{1}$ (resp. $S_{2}$) if it is a not maximal clique of $G_{1}$ (resp. $G_{2}$). Hence minimal separators of $G$ are single vertices and it follows that $G$ is a block graph.
Lemma 2.8.
@@
Proof 2.9.
From Lemma 2, $G$ is chordal and its minimal separators are the ones of $G_{1}$ and $G_{2}$ plus possibly $S_{1}$ or $S_{2}$ if they are not maximal cliques. If $S_{i}$ intersects exactly one maximal clique, $i\in\{1,2\}$, it is clear that it does not intersect any minimal separator of $G_{i}$. If $S_{i}$ is a critical clique, we claim that either $S_{i}$ is a minimal separator of $G_{i}$ or does not intersect any minimal separator of $G_{i}$. Indeed, if there is a minimal (clique) separator $S$ of $G_{i}$ such that $S_{i}\subset S$, then since $S_{i}$ is a critical clique, there exist $u\in S_{i},v\in S\setminus S_{i}$ and $w\in V(G_{i})\setminus S$ such that $w$ is adjacent to exactly one of $u$ and $v$. Suppose w.l.o.g. that $w$ is adjacent to $u$ and not to $v$, then $u$ is in a minimal $vw$-separator that intersects $S$ and is not equal to it. This is a contradiction since $G_{i}$ is strictly chordal and by Theorem 4 its minimal separators are pairwise disjoint. It follows that the minimal separators of $G$ are pairwise disjoint and by Theorem 4 $G$ is strictly chordal.
@@
Observation 1
Let $G_{1}=(V_{1},E_{1}),\dots,G_{r}=(V_{r},E_{r})$ be disjoint strictly chordal graphs and for each $i\in\{1,\dots,r\}$ let $S_{i}\subseteq V_{i}$ be a critical clique, a maximal clique or a set that intersects exactly one maximal clique of $G_{i}$. If $S_{1}$ is a critical clique or intersects exactly one maximal clique of $G_{1}$, then $S_{1}$ is a critical clique or intersects exactly one maximal clique of $(G_{1},S_{1})\otimes(G_{2},S_{2})$. This implies by Lemma 2.8 that $(G_{1},S_{1})\otimes(\bigcup_{2\leqslant i\leqslant r}G_{i},\bigcup_{2\leqslant i\leqslant r}S_{i})$ is strictly chordal (see Figure 3).
@@
We are left with Item iii, which can be proven in a similar way as the previous one. Using the same notation, the only vertices of $T_{A^{\prime}}$ whose degree may increase in $T_{A}$ are in $N$. Therefore, every connected component of $T_{A^{\prime}}\setminus(A^{\prime}\cup d(T_{A^{\prime}}))$ that does not contain vertices of $N$ is also a connected component of $T_{A}\setminus(A\cup d(T_{A}))$. Moreover, a connected component of $T_{A^{\prime}}\setminus(A^{\prime}\cup d(T_{A^{\prime}}))$ may contain vertices of $N$ only if these vertices have degree at most $2$ in $T_{A^{\prime}}$ and are not in $A$. In particular $|N|\leqslant 2$and both vertices lie in the same connected component of $T_{A^{\prime}}\setminus(A^{\prime}\cup d(T_{A^{\prime}}))$ (if any). Therefore, this connected component may be split in two disjoint ones in $T_{A}\setminus(A\cup d(T_{A}))$ if the vertices of $N$ are removed. Moreover, an additional connected component corresponding to the path between $a$ and $u$ may exist in $T_{A}$. Hence, we can only increase the number of connected components of $T_{A}\setminus(A\cup d(T_{A}))$ by compared to $T_{A^{\prime}}\setminus(A^{\prime}\cup d(T_{A^{\prime}}))$. Since the property is verified for $|A|=1$, it follows by induction that there are $2\cdot|A|$ connected components in $T_{A}\setminus(A\cup d(T_{A}))$.